exact dx+(x^2y+4y)dy=0 y(4)=0


1
Expert's answer
2021-10-12T10:17:39-0400

Let us solve the differential equation "dx+(x^2y+4y)dy=0,\ y(4)=0."

Since "\frac{\partial (1)}{\partial y}=0" but "\frac{\partial (x^2y+4y)}{\partial x}=2xy\ne 0," we conclude that this differential equation is not exact.

Let us solve it. For this let us divide both parts by "x^2+4." Then we have the equivalent differential equation "\frac{dx}{x^2+4}+ydy=0." It follows that "\int\frac{dx}{x^2+4}+\int ydy=C," and hence "\frac{1}2\arctan\frac{x}2+\frac{y^2}2=C."

Taking into account that "y(4)=0," we conclude that "C=\frac{1}2\arctan\frac{4} 2+0=\frac{1}2\arctan 2."

Therefore, "\arctan\frac{x}2+y^2=\arctan 2" is the solution of the differential equation "dx+(x^2y+4y)dy=0,\ y(4)=0."